OpenStudy (anonymous):

The volume of a square based pyramid is eighty-one cubic meters. What is the volume of the cube with the same side length and height?

OpenStudy (anonymous):

V = (1/3)b^2h where b = base length, h = height from vertex to the base V = 13 b^3 because base and height are the same Since the cube has volue of 81 m^3, then b^3 = 81 b = 81^(1/3) V = 13 ( 81^(1/3) ) V = (1/3)(81) V = 27 m^3

OpenStudy (anonymous):

V = 27 m^3 is the answer